OpenStudy (anonymous):

a cylinder of volume 1.22*10^-3 m^3 contains methane at 15 degrees centigrade. the pressure gauge on the cylinder shows the methane to be at a pressure of 3.28*10^5 Pa. what volume of O2 at 25 degrees centigrade and a pressure of 1*10^5 Pa will be required to completely burn all the methane? gas constant R=8.314JK^-1 mol^-1

OpenStudy (anonymous):

PV=nRT P=3.28*10^5 Pa V=1.22*10^-3 m^3 R=8.314JK^-1 mol^-1 T=15 degrees centigrade=(273+15)K=288K n=PV/RT=(3.28*10^5*1.22*10^-3)/(8.314*288) n=0.167moles CH4(g) + 2 O2(g) → CO2(g) + 2 H2O(l) ___________________________________________________________ 1mole of methane combines with 2 moles of O2 0.167 of methane combines with x moles of O2 1/2=0.167/x x=0.334moles of O2 ___________________________________________________________ P(oxy)=1*10^5 V(oxy)=? n=0.334moles T=25 degrees centigrade=(273+25)K=298K V(oxy)=nRT/P(oxy) V(oxy)=(0.334*8.314*298)/1*10^5 =0.00827509048= =8.275*10^-3cubic meters. volume of O2 required to completely burn all the methane=8.275*10^-3cubic meters.

OpenStudy (anonymous):

Thank you!!